OpenStudy (anonymous):

how can you solve tan{arctan[(x+1)/(x-1)] + arctan [(x-1)/x]}

OpenStudy (anonymous):

draw it.

OpenStudy (anonymous):

I tried letting a=(x+1)/(x-1) and b= (x-1)/x

OpenStudy (anonymous):

\[\tan [\tan^{-1}\frac{ x+1 }{ x-1 } +\tan^{-1} \frac{ x-1 }{ x }]\]

OpenStudy (anonymous):

 Hmmmm

OpenStudy (anonymous):



OpenStudy (anonymous):

I mean how do you solve it analytically?

OpenStudy (anonymous):

drawing is a good way to analyze.

OpenStudy (anonymous):

can I solve it as arctan a+arctanb = arctan a+b/1-ab ?

OpenStudy (anonymous):

Is that a property you are given?

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

you can also use angle addition property for tangent \[\tan(a+b) = \frac{\tan a +\tan b}{1 - \tan a \tan b}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

I got it now